OpenStudy (anonymous):

Solve for x: log3(x+4)+log3(x-2)=2

OpenStudy (zale101):

change the equation into the exponential form, so it can be easier to solve for x

OpenStudy (anonymous):

is it log base 3( x+1)or just 3*x+1

OpenStudy (zale101):

\[ \log_3(x+4)+\log_3(x-2)=2\] \[(x+4)(x-2)=2^3 \] make sense what i did here, on how i transformed the logarithm equation into the exponential form

OpenStudy (zale101):

@WAyne5

OpenStudy (zale101):

@nkaadi1995 3 is the base, i believe

OpenStudy (anonymous):

its zale 101 is correct except its 3^2 not 2^3

OpenStudy (zale101):

oh, and when u add logs together, u multiply, using the product rule

OpenStudy (anonymous):

yes

OpenStudy (zale101):

Solve for x: log3(x+4)+log3(x-2)=2 becomes log3(x+4)l(x-2)=2

OpenStudy (zale101):

oh yea lol @nkaadi1995

OpenStudy (anonymous):

no as you wrote before its multiply but the mistake is in 2^3

OpenStudy (zale101):